# ContextSeg: Sketch Semantic Segmentation by Querying the Context with Attention

### 28 Nov 2023 v1

@@ 1 Introduction
Sketches are widely used as a human-computer interaction tool. Many studies continuously explore its capabilities in modeling [12], retrieval [13], and generation [2]. Among these tasks, sketch interpretation serves as the foundation. However, it still remains challenging for computers due to the inherent ambiguity and sparsity of user sketches [28]. In this paper, we focus on sketch semantic segmentation, an essential task in finer-level sketch interpretation.
@@
Sketches are typically represented in three data formats, i.e., the raster image, graph, and point sequence. Recently, many learning-based approaches have adopted the aforementioned sketch representations revealing various advantages and disadvantages. Specifically, image-based methods [14, 32, 33] take raster images as input, exploiting absolute coordinates to capture the proximity between strokes. However, they tend to perform poorly in learning the structural information of strokes and overlook the sequential relationships between them. Graph-based methods [28, 26, 31] take graphs as input, which consists of stroke points as nodes and point connections as edges. The graph representation also exploits absolute coordinates, which facilitate proximity learning between strokes but is challenging to capture the structural information of a stroke. Sequence-based methods [9, 27, 19, 13] take point sequences as input, which uses relative coordinates to better encode the structural information of individual strokes. However, they struggle to capture the proximity relationships and spatial information between strokes. Nevertheless, the stroke structural and positional information and proximity relationship between strokes are essential for the sketch interpretation task, i.e., semantic segmentation in our case.
@@
Motivated by the observations, we propose ContextSeg, a two-stage approach marrying the advantages of all three representations. Specifically, at the stroke level, we treat an entire stroke instead of the stroke point as a single entity, represented as a raster image, where the positional information can be efficiently extracted by an autoencoder CNN network. Furthermore, due to the sparsity of stroke pixels, as well as the large geometric variation of a stroke depicting the drawing content, it is challenging to encode stroke structural information only by reconstruction. We thus propose a novel dense distance field regression task upon reconstruction, reinforcing structural information learning. At the sketch level, we treat all strokes within a sketch as a sequence of images. By augmenting the sequence with positional information, neural networks can well capture the proximity between strokes. As for the segmentation task, we exploit an auto-regressive Transformer, where instead of labeling one stroke at each time, we select a set of strokes belonging to the same semantic group at one time. Using the group-based prediction, context information, e.g., strokes possessing a part label and strokes remaining to be labeled, can be explicitly utilized by the Transformer decoder when making decisions for remaining strokes.
@@ 2 Related Work
Sketch Representation. In the field of computer vision, the representation learning of sketches has garnered widespread attention. As a fundamental task, its research success has been widely applied to multiple downstream tasks, i.e., sketch-based image retrieval (SBIR) [21], sketch generation [20], and sketch classification [29]. Zhang et al. [29] introduced deep convolutional neural networks (CNNs) for learning sketch embedding. The network took triplets composed of sketches, positive real images, and negative real images as input to discover coherent visual structures between the sketch and its positive pairs. Sain et al. [21] employed a cross-modal variational autoencoder (VAE) to learn sketch embedding. They explicitly decomposed each sketch into the semantic content shared with corresponding photos and the stylistic aspects unique to the sketch author. Furthermore, research has been conducted on the generalization properties of sketch embedding [18].
Semantic Sketch Segmentation. Compared to previous works that relied on handcrafted features and complex models [23, 22, 3], many deep learning-based models have achieved outstanding performance in sketch semantic segmentation tasks [28, 27, 31]. Based on the data formats they use, existing methods can be broadly grouped into three classes: image-based methods [14, 32, 33], sequence-based methods [27, 19, 13], and graph-based [28, 26, 31] methods. Zhu et al. [32] proposed an image-based dual convolutional neural networks (CNNs), where one branch with large-sized convolutional kernels was designed to handle long sketches, and another branch with small-sized convolutional kernels was tailored for processing short sketches. Li et al. [13] proposed a sequence-based encoder-decoder architecture for semantic sketch segmentation, applying additional constraints in the loss function, specifically targeting reconstruction and global grouping consistency. Yang et al. [28] proposed a graph-based network and utilized two graph convolutional branches to extract the inter-stroke features and the intra-stroke features.
@@
Vision Transformer. In addition to achieving tremendous success in the field of natural language processing [4], Transformer [24] is widely applied in numerous areas of computer vision, e.g., image classification [5, 6], object detection [1], segmentation [16], and point cloud analysis [8]. Transformer was initially used in auto-regression models in the field of natural language processing. However, many works have abandoned this strategy in computer vision tasks. Carion et al. [1] employed parallel decoding, given the absence of inherent order or sequence information between different bounding boxes in object detection tasks. Transformer was also adopted in the sketch domain, where Ribeiro et al. [20] encoded free-hand sketches in a vector format and effectively improved the performance of sketch-based image retrieval (SBIR) while Li et al. [12] treated sketch-based CAD modeling as a serializable translation problem and employed a Transformer-based network for stroke grouping.
@@ 3.1 Stroke Embedding
Embedding of a Group of Strokes. Other than the single strokes, we also encode a set of strokes corresponding to a semantic group, which is used in our segmentation Transformer. Specifically, suppose $g_{i}$ is a group of strokes, and $I(g_{i})$ is the corresponding image. We use all these group images as extra training data for the aforementioned embedding network training and obtain the group embedding $g_{i}^{e}$ as same as for a single stroke.
@@ 4 Experiments
Datasets. Following the common practice, we use both the SPG [13] and CreativeSketch [7] (dubbed as CS) datasets in this paper. Specifically, the SPG dataset contains sketches from 25 categories, each with 800 sketches. As in [13, 28], we use the same 20 selected categories and employ the same data splitting as described in [27], i.e., 700 sketches for training, 100 sketches for testing. We use all the training data from 20 categories for the embedding network training, while the segmentation Transformer is trained per category. As described in Sec. 3.1, we collected both single strokes and groups of strokes to train the embedding network. As for the CreativeSketch [7] dataset, it consists of 2 categories with much more complex sketches forming 8 and 17 semantic parts, respectively. We then randomly select 3000 sketches from each category with a 2500-500 training and testing split. Similarly, we use all 5000 sketches from both categories to train the embedding networks, while training the segmentation Transformer per category. All the sketches are scaled to fit into $256\times 256$ images.
Evaluation Metrics. We have adopted three metrics [10, 27, 12] to evaluate performance, i.e., stroke accuracy (SAcc), grouping accuracy (GAcc) and component accuracy (CAcc). Accuracy definitions can be found in the supplementary. Intuitively, stroke accuracy provides a fine-grained evaluation inspecting each stroke, while both the grouping accuracy and component accuracy quantify the semantic-based accuracy of all strokes. Compared with GAcc, CAcc is more vulnerable to the long tail problem regarding the stroke distribution over semantic groups.
@@
Ours w.o. CC 85.6 87.5 82.4 91.1 92.1 89.2 91.7 90.3 87.3 87.9 90.1 86.4 89.0 90.1 86.8
Ours w.o. DF 90.9 92.4 86.5 96.7 97.1 94.5 94.6 95.1 92.3 93.4 94.5 90.4 93.9 94.8 90.9
Ours-PD 91.1 92.5 86.8 97.2 97.8 95.1 95.3 96.2 93.1 94.1 95.3 91.2 94.4 95.4 91.6
@@ 4.1 Comparison
Quantitative comparisons are shown in Tab. 1, wherein we obtain the best average and per-category results (except for the face category) over all three metrics on both the SPG and CS datasets. Specifically, in terms of component accuracy, our approach achieved notable advancements of $1.4\%$, $1.8\%$, and $18.7\%$ compared to Sketch-Segformer, SketchGNN, and SPGSeg, respectively, on the SPG dataset. Also, on the more challenging CS dataset, substantial improvements in component accuracy were observed, reaching $11.9\%$, $13.3\%$, and $22.9\%$ in comparison to Sketch-Segformer, SketchGNN, and SPGSeg, respectively. Regarding the face category, our method demonstrates comparable accuracies, albeit slightly lower. This outcome is attributed to the presence of highly overlapped strokes, particularly representing hairs and eyes, which pose challenges to precise embedding learning and subsequently impact grouping predictions. In Fig. 4, qualitative results are showcased, illustrating our method’s capability to successfully identify strokes delineating intricate semantic details. Notably, it accurately captures elements such as clock numbers, as well as the wings and tail of a bird. More analysis and failure cases can be found in the supplementary.
@@ 4.2 Ablation Study
Parallel Decoding. Other than the used auto-regressive decoding, the parallel decoding scheme is frequently used in Transformer training and testing. Following [1], we thus change our auto-regressive decoding to the parallel decoding scheme (denoted as Ours-PD) and report the evaluation metrics in Tab. 2. The learned decoder input embeddings cannot explicitly access the context information, thus leading to notable performance degradation (e.g., 93.6% vs. 92.4% of the average CAcc).
@@ 5 Discussion
Table 3: The statistical comparison before and after the cross-category training.
@@
Total 94.8 96.1 +1.3 Total 95.6 95.9 +0.3 Total 97.8 98.3 +0.5 Total 98.8 99.0 +0.2 Total 93.2 93.9 +0.7
@@ 5.2 Sementic-aware Data Augmentation
Table 4: The statistical evaluation of each part before and after using the semantic-aware data augmentation.
@@ 6 Conclusion
In this paper, we proposed ContextSeg, for sketch semantic segmentation. We have achieved the best segmentation accuracy over SOTA methods. Comprehensive evaluations validate our superior performance, and preliminary experiments on cross-category training and semantic-aware data augmentation suggest inspiring research directions. However, our embedding network is not perfect and cannot encode strokes with rapid and large variations, while our segmentation Transformer fails to label strokes that are highly cluttered, e.g., the hairs and eyelashes.
@@ 7 More Results
7.1 Invariance Test.
@@ 7.2 Failure Cases
Figure 10 demonstrates several failure cases of our method. The imperfection primarily arises from two contributing factors. Firstly, our segmentation Transformer encounters difficulties in accurately labeling strokes within heavily cluttered areas. For instance, in Fig. 10 (a) and (b), distinguishing between hair and eyelash strokes becomes challenging due to their dense concentration, leading to segmentation inaccuracies. Similarly, in Fig. 10 (e) and (f), the intricate decorations on the basket pose challenges as our segmentation model erroneously categorizes them as part of the body. Secondly, our embedding network encounters limitations in encoding strokes characterized by rapid and substantial variations. For instance, in Fig. 10 (c) and (d), the leaves and ornamentation of the pineapple exhibit strokes with rapid fluctuations. Such variations can result in suboptimal embeddings, potentially causing misinterpretations for our segmentation Transformer. Consequently, certain strokes might be inaccurately classified as part of the body.

### 25 Mar 2024 v2

@@ 1 Introduction
Sketches are widely used as a human-computer interaction tool. Many studies keep exploring its capabilities in modeling [14, 13], retrieval [3, 2] and generation [5, 1]. Among these tasks, sketch interpretation serves as the foundation. However, it still remains challenging for computers due to the inherent ambiguity and sparsity of user sketches [31]. In this paper, we focus on sketch semantic segmentation, an essential task in finer-level sketch interpretation.
@@
Sketches are typically represented in three data formats, i.e., the raster image, graph and point sequence. Recently, many learning-based approaches have adopted the aforementioned sketch representations revealing various advantages and disadvantages (see the inset table). Specifically, image-based methods [16, 35, 36] take raster images as input, exploiting absolute coordinates to capture the
[…]
proximity between strokes. However, they tend to perform poorly in learning the structural information about strokes and overlook the sequential relationships between them. Graph-based methods [31, 29, 34] take graphs as input, which consists of stroke points as nodes and point connections as edges. The graph representation also exploits absolute coordinates, which facilitate proximity learning between strokes but is challenging to capture the structural information of a stroke. Sequence-based methods [10, 30, 22, 15] take as input point sequences, which use relative coordinates to better encode structural information of individual strokes. However, they struggle to capture the proximity relationships and spatial information between strokes. Nevertheless, stroke structural and positional information and the proximity relationship between strokes are essential for sketch interpretation.
@@
Motivated by the observations, we propose ContextSeg, a two-stage approach marrying advantages of all three representations to capture better structural and positional information of each single stroke, and the proximity relationship between strokes. Specifically, at the stroke level, we treat an entire stroke instead of the stroke point as a single entity, represented as a raster image, where the positional information can be efficiently extracted by an autoencoder CNN network. Furthermore, due to the sparsity of stroke pixels, as well as the large geometric variation of a stroke depicting the drawing content, it is challenging to encode stroke structural information only by reconstruction. We thus propose a novel dense distance field regression task upon reconstruction, reinforcing structural information learning. At the sketch level, we treat all strokes within a sketch as a sequence of images. By augmenting the sequence with positional information, neural networks can well capture the proximity between strokes. As for the segmentation task, we exploit an auto-regressive Transformer, where instead of labeling one stroke at each time, we select a set of strokes belonging to the same semantic group at one time. Using the group-based prediction, context information, e.g., strokes possessing a part label and strokes remaining to be labeled, can be explicitly utilized by the Transformer decoder when making decisions for remaining strokes.
@@ 2 Related Work
Sketch Representation. In the field of computer vision, sketch representation learning has garnered widespread attention. It is a fundamental task in numerous downstream applications, i.e., sketch-based image retrieval (SBIR) [24, 3], sketch generation [23], and sketch classification [32]. Zhang et al. [32] introduced deep convolutional neural networks for learning sketch embedding. The network took triplets composed of sketches, positive real images, and negative real images as input to discover coherent visual structures between the sketch and its positive pairs. Sain et al. [24] employed a cross-modal VAE to disentangle sketches into shared semantic content and unique style contents, enhanced with meta-training for dynamic adaptation to unseen user styles, enabling style-agnostic SBIR. Furthermore, research has been conducted on the generalization properties of sketch embedding [21, 17, 1, 9].
Semantic Sketch Segmentation. Compared to previous works that relied on handcrafted features and complex models [26, 25, 6], many deep learning-based models have achieved outstanding performance in sketch semantic segmentation tasks [31, 30, 34]. Based on the data formats they use, existing methods can be broadly grouped into three classes: image-based [16, 35, 36], sequence-based [30, 22, 15], and graph-based [31, 29, 34] methods. Zhu et al. [35] proposed a dual-CNN approach for sketch segmentation and labeling, employing two networks with distinct kernel sizes to handle different sketch lengths, while enhancing performance through the integration of position and orientation as a triple-channel input using fused masks. Li et al. [15] proposed a sequence-based encoder-decoder architecture for sketch semantic segmentation, applying additional constraints in the loss function, specifically targeting reconstruction and global grouping consistency. Yang et al. [31] proposed a graph-based network and utilized two graph convolutional branches to extract the inter-stroke features and the intra-stroke features.
@@
Vision Transformer. Other than natural language processing, Transformer [27] has been widely adopted in all kinds of computer vision tasks, e.g., classification [7, 8], detection [4], segmentation [19]. Unlike the auto-regressive decoding strategy, Carion et al. [4] proposed parallel decoding, given the absence of inherent order or sequence information between different bounding boxes in object detection tasks. Transformer was also adopted in the sketch domain, where Ribeiro et al. [23] encoded free-hand sketches in a vector format and effectively improved the performance of sketch-based image retrieval while Li et al. [14] treated sketch-based CAD modeling as a serializable translation problem and employed a Transformer-based network for stroke grouping.
@@ 3.1 Stroke Embedding
Embedding of a Group of Strokes. Other than the single strokes, we also encode a set of strokes corresponding to a semantic group, which is used in our segmentation Transformer. Specifically, suppose $g_{i}$ is a group of strokes, and $I(g_{i})$ is the corresponding image. We use all these group images as extra training data for the aforementioned embedding network training and obtain the group embedding $g_{i}^{e}=E^{s}(I(g_{i}))$ as same as for a single stroke.
@@ 4 Experiments
Datasets. Following existing work, we use both the SPG [15] and CreativeSketch [9] (dubbed as CS) datasets in this paper. Specifically, the SPG dataset contains sketches from 25 categories, each with 800 sketches. As in [15, 31], we use the same 20 selected categories and employ the same data splitting as described in [30], i.e., 700 sketches for training, 100 sketches for testing. We use all the training data from 20 categories for the embedding network training, while the segmentation Transformer is trained per category. As described in Sec. 3.1, we collected both single strokes and groups of strokes to train the embedding network. As for the CS dataset, it consists of 2 categories with much more complex sketches forming 8 and 17 semantic parts, respectively. We then randomly select 3000 sketches from each category with a 2500-500 training and testing split. Similarly, we use all 5000 sketches from both categories to train the embedding networks, while training the segmentation Transformer per category. All the sketches are scaled to fit into $256\times 256$ images.
Evaluation Metrics. We have adopted three metrics [11, 30, 14] to evaluate performance, i.e., stroke accuracy (SAcc), grouping accuracy (GAcc) and component accuracy (CAcc), and their definitions can be found in the supplementary. Intuitively, stroke accuracy provides a fine-grained evaluation inspecting each stroke, while both the grouping accuracy and component accuracy quantify the semantic-based accuracy of all strokes. Compared with GAcc, CAcc is more vulnerable to the long tail problem regarding the stroke distribution over semantic groups.
@@
Ours w/o CC 85.6 87.5 82.4 91.1 92.1 89.2 91.7 90.3 87.3 87.9 90.1 86.4 89.1 90.0 86.3
Ours w/o DF 90.9 92.4 86.5 96.7 97.1 94.5 94.6 95.1 92.3 93.4 94.5 90.4 93.9 94.8 90.9
Ours-PD 91.1 92.5 86.8 97.2 97.8 95.1 95.3 96.2 93.1 94.1 95.3 91.2 94.4 95.5 91.6
Ours 93.2 94.9 89.5 99.2 98.7 97.5 96.4 98.4 94.8 95.9 96.5 92.4 96.2 97.1 93.6
@@ 4.1 Comparison
Quantitative comparisons are shown in Tab. 1, wherein we obtain the best average and per-category results (except for the face category) over all three metrics on both the SPG and CS datasets. Specifically, in terms of component accuracy, our approach achieved notable advancements of $1.4\%$, $1.7\%$, and $18.7\%$ compared to Sketch-Segformer, SketchGNN, and SPGSeg, respectively, on the SPG dataset. Also, on the more challenging CS dataset, substantial improvements in component accuracy were observed, reaching $11.9\%$, $13.3\%$, and $22.9\%$ in comparison to Sketch-Segformer, SketchGNN, and SPGSeg, respectively. Regarding the face category, our method demonstrates comparable accuracies, albeit slightly lower. This outcome is attributed to the presence of highly overlapped strokes, particularly representing hairs and eyes, which pose challenges to precise embedding learning and subsequently impact grouping predictions. In Fig. 4, qualitative results are showcased, illustrating our method’s capability to successfully identify strokes delineating intricate semantic details. Notably, it accurately captures elements such as clock numbers, as well as the wings and tail of a bird. More analysis and failure cases can be found in the supplementary.
@@ 4.2 Ablation Study
Parallel Decoding. Other than the used auto-regressive decoding, the parallel decoding scheme is frequently used in Transformer training and testing. Following [4], we thus change our auto-regressive decoding to the parallel decoding scheme (denoted as Ours-PD) and report the evaluation metrics in Tab. 2. The learned decoder input embeddings cannot explicitly access the context information, thus leading to notable performance degradation (e.g., 93.6% vs. 91.6% of the average CAcc).
Additional Ablation Study. We have conducted more experiments on a) positional encoding of strokes and b) distance field-only decoding in the embedding learning, c) stroke-based auto-regressive decoding, and d) the group order in auto-regressive decoding, more results statistical and visual results can be found from the supplementary.
@@ 5 Discussion
Table 3: The statistical comparison before and after the cross-category training on SAcc.
@@
Average 88.4 92.0 +3.6 Average 94.2 94.4 +0.2 Average 96.5 97.5 +1.0 Average 98.8 99.0 +0.2 Average 90.9 91.0 +0.1
@@ 5.2 Sementic-aware Data Augmentation
Table 4: The statistical evaluation of each part before and after using the semantic-aware data augmentation on SAcc.
@@ 6 Conclusion
In this paper, we proposed ContextSeg, for sketch semantic segmentation, achieving the best segmentation accuracy over SOTA methods. Comprehensive evaluations validate our superior performance, and preliminary experiments on cross-category training and semantic-aware data augmentation suggest inspiring research directions. However, our embedding network is not perfect and cannot encode strokes with rapid and large variations, while our segmentation Transformer fails to label strokes that are highly cluttered, e.g., the hairs and eyelashes.
@@ 7 More Results
7.1 Additional Ablation Study
Positional Encoding. In stroke embedding learning (Sec. 3.1), we have used two additional coordinate channels to augment the stroke information. To understand better its effectiveness, we replaced the 2D coordinates with the popular 2D sinusoidal positional encoding (denoted as 2DPE and Ours-PE for embedding and segmentation). The stroke reconstruction result and the evaluation statistics are shown in Fig. 10 and Tab. 5, respectively, where the reconstructed sketch is blurry, and the segmentation metrics are all significantly inferior to ours.
Distance Field-only Embedding. To further validate the efficacy of the distance field prediction branch, we have trained the embedding network with only the distance field (denoted as DF). The predicted distance field is shown in Fig. 11, where individual strokes are barely recognized. Besides, we have linked the embedding network with our segmentor (denoted as Ours-DF), and report the evaluation metrics in Tab. 5. The results are inferior to Ours and even worse than Ours w/o DF because the sketch instead of a dense and rough approximation is the key to the segmentation task.
Stroke-based Decoding. The design philosophy of the group-based prediction and its effectiveness are discussed in Sec. 3.2 and Sec. 4.2. We further experiment with stroke-based auto-regressive decoding (denoted as Ours-S) since it is more intuitive. Statistics are shown in the Tab. 5, where our method achieved a remarkable $4.4\%$ increase in terms of average SAcc. Besides, group-based prediction is more efficient, e.g., the average inference time on the airplane category is two times faster (0.73s vs. 1.86s).
Group Order in Auto-regressive Decoding. By design, the stroke order serving as the positional encoding in the Transformer encoder does not matter the decoder prediction, however, the group order matters. By default, we empirically use the more intuitive stroke frequency-based descend order. We have tried two alternatives: stroke frequency ascend order (denoted as Ours-InvG), and a random order (denoted as Ours-RanG). The statistics are shown in Tab. 5, where ours archives the best while the other two are on par with ours.
Figure 10: Sketch reconstruction from 2DPE and Ours.
Figure 11: Distance field prediction from DF.
Table 5: Statistical results of additional ablation studies on the SPG dataset.
Method Airplane Calculator Face Ice cream Average
Ours-PE 89.3 91.9 85.6 96.1 95.7 96.8 93.4 95.4 93.4 93.1 94.5 91.2 93.0 94.4 91.8
Ours-DF 83.4 86.6 81.3 90.1 91.7 87.2 89.4 89.2 86.3 86.7 89.3 84.2 87.4 89.2 84.8
Ours-S 88.6 91,1 84.7 94.4 95.8 93.6 92.1 94.6 91.7 91.8 92.3 89.6 91.7 93.5 89.9
Ours-InvG 92.6 94.0 88.6 98.6 98.1 97.0 95.9 97.9 93.8 95.2 96.1 91.7 95,6 96.5 92.8
Ours-RanG 92.8 94.1 88.7 98.7 98.2 97.1 96.1 98.0 94.1 95.4 96.2 91.9 95.8 96.8 93.0
7.2 Invariance Test
@@ 7.3 Failure Cases
Figure 12 demonstrates several failure cases of our method. The imperfection primarily arises from two contributing factors. Firstly, our segmentation Transformer encounters difficulties in accurately labeling strokes within heavily overlapped areas. For instance, in Fig. 12 (a) and (b), distinguishing between hair and eyelash strokes becomes challenging due to their dense concentration, leading to segmentation inaccuracies. Similarly, in Fig. 12 (e) and (f), the intricate decorations on the basket pose challenges as our segmentation model erroneously categorizes them as part of the body. Secondly, our embedding network encounters limitations in encoding strokes characterized by rapid and substantial variations. For instance, in Fig. 12 (c) and (d), the leaves and ornamentation of the pineapple exhibit strokes with rapid fluctuations. Such variations can result in suboptimal embeddings, potentially causing misinterpretations for our segmentation Transformer. Consequently, certain strokes might be inaccurately classified as part of the body.